OpenStudy (anonymous):

Solve each system of linear equations using matrices. x+_y=1 x-2y=4

OpenStudy (zzr0ck3r):

what os _y?

OpenStudy (anonymous):

theres a space there

OpenStudy (zzr0ck3r):

? x+y = 1 x-2y = 4 ?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

Ok, so the matrix would looke like [1,1,1;1, -2, 4] correct? where [first row;second row]

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

\[\begin{bmatrix} 1& 1\\ 1& -2 \end{bmatrix} \begin{bmatrix} x\\ y \end{bmatrix}= \begin{bmatrix} 1\\ 4 \end{bmatrix}\]

OpenStudy (zzr0ck3r):

what do you need to add to r1c2 to get it to be 0?

OpenStudy (anonymous):

I don't know that's why I'm asking for help

OpenStudy (zzr0ck3r):

i asked what you need to add to 1 to get 0

OpenStudy (anonymous):

\[A\begin{bmatrix} x\\ y \end{bmatrix}= \begin{bmatrix} 1\\ 4 \end{bmatrix}\] \[A^{-1}A\begin{bmatrix} x\\ y \end{bmatrix}=\begin{bmatrix} x\\ y \end{bmatrix}= A^{-1}\begin{bmatrix} 1\\ 4 \end{bmatrix}\]

OpenStudy (zzr0ck3r):

this is about row reduction henpen

OpenStudy (anonymous):

Basically, just like in normal algebra, we 'divide by A' (or multiply by the inverse matrix of a) to get what x and y actually ARE.

OpenStudy (zzr0ck3r):

I doupt they have got to inverse yet

OpenStudy (anonymous):

I'm so confused

OpenStudy (zzr0ck3r):

are you doing row reduction in class?

OpenStudy (anonymous):

I think so

OpenStudy (zzr0ck3r):

have you been to class?

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

ok would you be able to solve this with elelmination method we learn in algebra class?

OpenStudy (anonymous):

I'm sorry that math isn't an easy subject for me. Thanks for making it a joke though. Nice. I'll just find someone else to explain it to me.

OpenStudy (zzr0ck3r):

I did not make a joke, I asked what you are doing in class. You need not understnad exactly how to do what you are doing in class, but you should know what it is you are doing.

OpenStudy (zzr0ck3r):

if you want help. there are many ways to solve such a question....so knowing what you are working on could be very helpfull.

OpenStudy (zzr0ck3r):

but ok good luck m8

OpenStudy (zzr0ck3r):

what you say we start over?:) would you be able to solve this with the elimitation method we learn in algebra @mluna1218 ?

OpenStudy (anonymous):

I figured this out by myself. Thanks though